Solve each of the following equations and check your solution by substituting in the equation.

(i) x/2-10=1/2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

x/2 – 10 = 1/2

(x – 20)/2 = 1/2

x – 20 = 1

x = 21

Thanks